IR SPECTROSCOPY STUDY ON THE (HCl)_{n}(H_{2}O)_{m} AGGREGATION IN HELIUM NANODROPLETS

\underline{PABLO NIETO}, MELANIE LETZNER, DANIEL HABIG, TOERSTEN POERSCHKE, SARAH ANGELIQUE GRÜN, KENNY HANKE, GERHARD SCHWAAB and MARTINA HAVENITH, Department of Physical Chemistry II, Ruhr-Universität Bochum, Germany.

The study of acid-water clusters is an active area of research due to its fundamental importance for chemistry^{a}^{,}^{b}. In particular the (HCl)_{n}(H_{2}O)_{m} clusters have been extensively investigated both theoretically and experimentally as a benchmark system. Despite of the great effort devoted to its understanding HCl dissociation in water clusters is still not well understood. An IR-Spectroscopy study on (HCl)_{n}(H_{2}O)_{m} embedded in helium nanodroplets will be presented. The H_{2}^{16}O→H_{2}^{18}O and isotopic substitution was used in the experiments to probe the bands in the 2650-2760 cm^{−1} spectral range which has been object of some debate recently^{c}^{,}^{d}. The observed isotopic shifts for the different bands raise some new questions to be addressed.

aD. Marx, Chem. Phys. Chem. 7, 1848, (2006).

bV. E. Bondybey et al., Int. Rev. Phys. Chem. 21, 277 (2002).

cA. Gutberlet et al., Science 324, 1545 (2009).

dS. D. Flynn et al., Phys. Chem. Lett. 1, 2233 (2010).